# Semantics-Preserving Evasion of LLM Vulnerability Detectors

### v1

@@ 5.4 Transfer Attacks Succeed on Black-Box APIs; On-Target Reveals Complex Interactions
The output format injection baseline achieves high success (55–99% across models), revealing that models can be manipulated by embedding target JSON outputs in code comments, an orthogonal vulnerability distinct from semantic robustness. In contrast, random_idsub is significantly weaker (7–30%), demonstrating that non-optimized renaming has a limited impact on code-understanding robustness.
@@ 6 Discussion
Clean-set accuracy alone does not guarantee reliability. While models like CodeAstra achieve near-70% clean recall, our evaluation reveals that this capability degrades under adversarial transformation, with complete resistance dropping to near zero (0.12%), implying that current detectors act less as robust semantic parsers and more as pattern-matchers over fragile lexical cues. Across semantics-preserving edits (§3), we observe a persistent integrity gap: models with high clean coverage can exhibit high union evasion and near-zero complete resistance, meaning that an attacker who can choose among multiple perturbation surfaces can often overturn most clean true positives. Even simple carrier-elimination preprocessing (§5.3) does not eliminate this risk, as evasion can migrate to remaining executable-looking channels.
Multiple metrics matter. Single-number summaries can be misleading for deployment decisions. Union success captures whether any available surface enables evasion, while complete resistance operationalizes a security floor when attackers can try multiple carriers. Per-method and CWE-level slices then identify where robustness breaks (which channels, which vulnerability types), enabling targeted mitigation rather than blanket conclusions from aggregate ASR.
Temperature and transfer analysis. The observed fragility is structural rather than stochastic. Evasion rates remain stable across decoding temperatures (Appendix D.2), ruling out random sampling artifacts. More importantly, the high transferability of attacks from a 14B surrogate model to GPT-4o points to shared representational blind spots: distinct models appear to over-index on identical non-executable tokens (e.g., #ifdef guards) while ignoring the invariant executable logic.
Decoupling of Domain Capability and Robustness. The performance profile of CodeAstra reveals a meaningful insight: although task-specific fine-tuning boosts baseline detection, it does not guarantee robustness. CodeAstra exhibits high vulnerability in specific channels (e.g., macro insertion) despite high clean accuracy. This suggests that domain specialization can coexist with-or mask-structural fragility, emphasizing that fine-tuned models require integrity evaluations just as rigorous as general-purpose ones.
Implications for Deployment and Defense. Our findings suggest three concrete directions for hardening detectors. First, training on carrier-diverse adversarial examples (spanning identifier, comment, and preprocessor surfaces) may enforce broader invariance than single-method defenses. Second, Complete Resistance should join clean accuracy as a co-equal deployment metric- a detector is only as robust as its weakest carrier. Third, ensemble approaches combining sanitization (for boundary carriers) with robust base models may offer defense-in-depth where neither alone suffices.
Limitations and Future Directions. We focus on C/C++ functions with a fixed prompting protocol; extending to other languages and detection formulations would strengthen generality. While Tier 2 validation (differential AST analysis) scales to 5,000 functions and Tier 3 (compilation-based) provides strong proof on 1,434 functions (Appendix A.3), absolute behavioral equivalence requires full execution environments unavailable in function-level datasets. Future work should explore adversarial fine-tuning on semantics-preserving carriers, though computational cost and overfitting risks motivate investigating lightweight defense wrappers (Madry et al., 2019; Varshney et al., 2024).
@@ 7 Conclusion
On a unified C/C++ benchmark (N=5000N{=}5000), we present an integrity evaluation of LLM-based vulnerability detectors under semantics-preserving code perturbations, evaluating five carriers spanning four insertion families (identifier substitution, comments, preprocessor insertions, and dead-branch insertion) and quantifying robustness via union success and complete resistance. Across models, we observe widespread failures of semantic invariance: many vulnerabilities that are correctly detected on clean inputs can be flipped to false negatives via behavior-equivalent edits, and a single universal, carrier-constrained string optimized on a surrogate can transfer unchanged to black-box API targets. Sanitization can remove some insertion surfaces, but it rewrites many inputs and leaves executable-looking channels vulnerable. These findings collectively suggest that clean benchmark performance alone is insufficient as a security guarantee, and that integrity-oriented evaluations and defenses should explicitly target robustness to behavior-preserving code changes without resorting to overly conservative task-altering preprocessing.
Impact Statement
This work evaluates the robustness of LLM-based vulnerability detectors under semantics-preserving adversarial perturbations. We discuss both the positive contributions and potential risks of this research.
Positive Impacts. Our findings reveal essential blind spots in current LLM-based code security tools before they are exploited in the wild. By systematically characterizing failure modes and introducing metrics such as Complete Resistance, we provide actionable diagnostics for practitioners to harden detection pipelines. We believe responsible disclosure of these weaknesses, consistent with standard practices in security research, enables the community to develop more robust defenses.
Potential Risks and Mitigations. As with most security research, the techniques described could theoretically be misused to evade vulnerability detection. However, we note that: (1) the attack surfaces we study (identifier renaming, comment insertion, preprocessor directives) are already well-known code transformations; (2) we do not release optimized adversarial strings or attack toolkits; and (3) the primary barrier to exploiting vulnerabilities lies in discovering them, not in evading detection. We believe the defensive value of exposing these weaknesses substantially outweighs the marginal risk of informing attackers about evasion strategies they could likely discover independently.
@@ A.1 Target Models Details
We evaluate the proposed attacks across five state-of-the-art LLMs, categorized by their access regimes:
@@
White-box CodeAstra rootxhacker/CodeAstra-7B (Ganesan, 2024)
@@
Operational Black-box Qwen2.5-Coder Qwen/Qwen2.5-Coder-32B-Instruct
@@ A.2 Vocabulary Constraints
A.3 Semantics Preservation Validation
We validate semantics preservation through a two-tiered approach: (1) Constraint-based Validation to verify method-specific injection rules, and (2) Differential Syntactic Analysis using Tree-sitter to ensure that perturbations do not introduce new syntax errors compared to the original code. This approach is robust to the incomplete nature of function-level datasets (e.g., missing headers) while strictly flagging syntax violations.
@@ A.3 Semantics Preservation Validation
Table 6: Semantics Preservation Rates across Models. We report two metrics: (1) Constraint Sat. Fraction of samples satisfying method-specific rules. (2) Tree-sitter Valid. Fraction of samples where the attack introduces no new AST error nodes ($E_{adv}\leq E_{orig}$). High validity across diverse architectures confirms the stability of our attack generation.
@@
    Evaluate on CodeAstra (highest TPR=69.08%) → 1,434 correctly detected (71.7%)
@@
Validation Results. Table 7 shows compilation preservation rates. Additive carriers achieve perfect 100% preservation (1,434/1,434), providing compiler-verified proof of semantics-preservation by construction. Identifier Substitution achieves 90.9% (1,303/1,434). The 131 failures stem from collisions with injected header symbols (uint8_t, size_t, printf, malloc, etc.) and are excluded from ASR calculations.
@@ B.1 Dataset Construction Details
Data sources and target language. Unified-VUL-N contains vulnerable functions written in C/C++ and isconstructed by pooling the vulnerable subsets from PrimeVul, BigVul, and DiverseVul.The goal is to evaluate robustness under semantics-preserving transformations on asingle, fixed benchmark with controlled composition.
@@
Context budget and token-length filtering. To guarantee sufficient headroom for injected strings and to standardize the context budget across models, we filter functions by code-only token length using the Qwen2.5-Coder-14B tokenizer. Specifically, we require the code-only length to be at most 4096 tokens, where “code-only” refers to the function text used as input code (excluding any additional prompt or JSON schema wrappers). This constraint ensures that our injected variants remain within model context limits under the shared evaluation prompt.
@@ B.3 Benign-Class Invariance and False Positives
Setup. We evaluate benign-class invariance on a held-out set of $N_{\mathrm{benign}}{=}500$ ground-truth benign functions. Clean FPR is the fraction of benign functions predicted as VULNERABLE under clean inference (baseline model error). Sanitized FPR is computed after applying the same sanitization pipeline as in §5.3 (comment stripping and preprocessor removal); we report $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ as the change in false-positive rate (percentage points) from clean to sanitized inputs. Benign-class Invariance Rate is the fraction of benign functions that are predicted as BENIGN under clean inference and remain BENIGN after applying our semantics-preserving attacks (union over all attack methods). This metric directly measures whether our attacks preserve benign-class predictions, which is aligned with our attack objective (evading vulnerability detection without introducing false positives).
Results. Table 8 reports per-model clean FPR, sanitized FPR (and $\Delta\mathrm{FPR}_{\mathrm{sanit}}$), and benign-class invariance rate. All models achieve 100% benign-class invariance rate, demonstrating that our semantics-preserving attacks preserve benign-class predictions: all functions that are correctly classified as BENIGN under clean inference remain BENIGN after attack.
Table 8: Benign-side effects and invariance. Clean/Sanitized FPRs are measured on $N_{\mathrm{benign}}$ ground-truth benign functions. $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ is the change (percentage points) after sanitization. Benign-class Invariance is computed on benign functions predicted as BENIGN under clean inference and remains BENIGN after applying the union of semantics-preserving attacks.
@@ B.4 Robustness of Reference Tokenizer for Length Filtering
To ensure fair cross-model comparisons, we standardized length filtering using the Qwen2.5-Coder-14B tokenizer to maintain an identical evaluation set across all targets. Per-model filtering would result in inconsistent subsets, making metrics like $TPR_{clean}$ and $ASR_{cond}$ non-comparable. Correlation analysis on 500 random samples confirms that the Qwen tokenizer is highly representative: it achieves a Pearson correlation of $r=0.9998$ with Llama-3.1-8B and $r=0.9941$ with Mistral-7B-Instruct-v0.2, which is the original model of CodeAstra. The mean length difference relative to Llama is only 5.57-5.57 tokens (max 16). Given our 4096-token limit, these negligible deviations ensure that the filtering process does not introduce systemic bias against specific architectures.
@@ B.5 Compute and Latency Statistics
Cost metrics. We track (i) forward passes for gradients ($m_{c}$ per step during curriculum, $m_{\mathrm{train}}$ per step after balance), (ii) forward passes for candidate evaluation ($m_{c}\cdot\lceil\texttt{search\_width}/\texttt{batch\_size}\rceil=m_{c}\cdot 8$ per step with our default settings), and (iii) wall-clock time (total optimization time including GPU memory transfers and batched operations). Optional per-step model.generate() calls are counted separately and can be disabled via enable_step_generate=False to reduce latency.
@@ C.1 Site Selection and Functional Role
Positional Sensitivity. Table 9 compares the $ASR_{cond}$ for different comment-carrier insertion sites. We observe that Trailing comments (placed after the function body) often exhibit high effectiveness, notably for Llama3.1-8B and CodeAstra, where success rates significantly exceed those of the original Leading header comments. However, Middle placements (interleaved with the logic) often show lower effectiveness across all models. This suggests that the model’s contextual focus on interleaved logic may partially mitigate the influence of semantics-preserving perturbations compared to injections at the function boundaries.
@@
Table 9: Ablation results for different injection sites and functional roles. $ASR_{cond}$ is reported as the fraction of originally correctly detected vulnerable functions ($TP_{clean}$) that were flipped to BENIGN.
Model Family Variant Flipped / $TP_{clean}$ $ASR_{cond}$ (%)
@@
Appendix D Extended Sanitization Analysis
@@
As shown in Table 10, the detection boundary is highly unstable. For Llama3.1-8B, simply removing developer comments (which have no effect on program execution) causes the model to ”discover” 133 previously missed vulnerabilities—an 8.6% increase. This suggests that the model’s clean accuracy is partly contingent on specific lexical distributions rather than robust semantic understanding. Qwen2.5-Coder-32B shows similar behavior (+5.2%), while CodeAstra exhibits minimal drift (+0.09%), suggesting different sensitivities to non-executable context across architectures.

### v2

@@ 5.5 Non-GCG Baselines
The output format injection baseline achieves high success (55–99% across models), revealing that models can be manipulated by embedding target JSON outputs in code comments, an orthogonal vulnerability distinct from semantic robustness. In contrast, random_idsub is substantially weaker (7–30%), confirming that non-optimized renaming has limited impact on code-understanding robustness.
For insertion-based carriers, we evaluated a random token insertion baseline: vocabulary-sampled tokens at equal string length, injected via the comment carrier with no gradient optimization. This achieves 13.2% $\mathrm{ASR}_{\mathrm{cond}}$ versus GCG’s 91.5% under identical conditions. The gap confirms that the detection boundary in token space is highly non-linear: random search in the unconstrained comment space rarely locates evasion-relevant directions, while gradient guidance does so reliably.
@@ 6 Discussion
Clean recall as a misleading proxy. Under the five syntax-preserving carriers evaluated, even high-accuracy detectors exhibit near-zero Complete Resistance, indicating reliance on non-executable tokens rather than program semantics. Union ASR captures whether any carrier enables evasion, while CR operationalizes a security floor across carriers. We note that several models (CodeAstra: 70% FPR, GPT-5-mini: 54% FPR on benign inputs; Appendix B.3) are questionable as standalone deployment gates even before adversarial pressure. Our CR metric complements broader multi-objective evaluation frameworks such as SecLens-R (Halder et al., 2026), adding a robustness dimension to stakeholder-aware vulnerability detection assessment.
Defense landscape. No evaluated defense is sufficient in isolation (§5.3). Input sanitization eliminates comment and preprocessor carriers but induces prediction drift of up to 8.6% and leaves identifier substitution and dead-branch carriers intact. Randomized smoothing yields only marginal reductions (1.6–8.9 pp). Adversarial fine-tuning provides the largest CR improvement (0.13% $\to$ 3.88%) but shifts vulnerability to the identifier carrier and reduces clean TPR by 6.3 pp. These results suggest that effective mitigation requires combining complementary defenses across inference and training time; carrier-diverse adversarial training (Madry et al., 2019) with broader carrier coverage and purification wrappers (Mu et al., 2025) are concrete next steps.
Limitations. We focus on C/C++ with a fixed prompting protocol; extension to other languages would strengthen generality. Our three-tier validation certifies syntax- and compilation-level preservation only; formal runtime equivalence requires execution environments unavailable in function-level datasets (Appendix A.3). The optimized universal strings (particularly for identifier substitution) are often unnatural and may be flagged by style linters or repository naming policies; the attack success rates reported here assume no such pre-screening. Multi-seed analysis covers on-target identifier substitution across three models; broader carrier and regime coverage remains future work.
@@ 7 Conclusion
On a unified C/C++ benchmark (N=5000N{=}5000), we evaluate six LLM-based vulnerability detectors under five syntax-preserving (Tree-sitter-validated) and largely compilation-preserving (gcc-validated) carrier variants. Complete Resistance falls to at most 13% for five of six models, and a universal string optimized on a 14B surrogate transfers unchanged to black-box APIs including GPT-4o (87% union $\mathrm{ASR}_{\mathrm{cond}}$). Among defenses, sanitization addresses only comment and preprocessor surfaces, and randomized smoothing yields marginal reductions (1.6–8.9 pp). Adversarial fine-tuning improves CR from 0.13% to 3.88% but introduces a robustness-accuracy trade-off. Multi-seed analysis across three models confirms that the observed exploitability is systematic rather than seed-dependent. These results indicate that clean benchmark recall is not a proxy for deployment robustness; integrity-oriented evaluations and training-time defenses must target robustness to syntax-preserving changes.
@@ A.1 Target Models Details
We evaluate the proposed attacks across six target LLMs and one surrogate, categorized by their access regimes:
@@
White-box CodeAstra rootxhacker/CodeAstra-7B Ganesan (2024)
White-box StarCoder2 bigcode/starcoder2-15b-instruct-v0.1
@@
Cost-limited Gray-box Qwen2.5-Coder Qwen/Qwen2.5-Coder-32B-Instruct
@@ A.2 Vocabulary Constraints
  • Comment carrier. We restrict to printable ASCII excluding sequences that could terminate the comment block (*/, newlines).
@@
A.3 Syntax Preservation and Compilation Validation
We validate syntax and compilation preservation through a three-tiered approach: (1) Constraint-based Validation to verify method-specific injection rules, (2) Differential Syntactic Analysis using Tree-sitter to ensure perturbations do not introduce new syntax errors compared to the original code (applied on all 5,000 functions), and (3) Compilation-based Validation via gcc on a 1,434-function compilable subset. All three tiers certify syntactic and compilation-level preservation only; formal runtime-equivalence would require full execution environments unavailable in function-level datasets.
@@ A.3 Syntax Preservation and Compilation Validation
Table 3: Syntax Preservation Rates across Models (Tree-sitter Tier). We report two metrics: (1) Constraint Sat. Fraction of samples satisfying method-specific rules. (2) Tree-sitter Valid. Fraction of samples where the attack introduces no new AST error nodes ($E_{adv}\leq E_{orig}$). These rates certify syntactic preservation only; compilation-level validation is reported separately in Table 4. High validity across diverse architectures confirms the stability of our attack generation.
@@
    Evaluate on CodeAstra (TPR=69.08%; primary white-box target) $\rightarrow$ 1,434 correctly detected (71.7%)
@@
Validation Results. Table 4 shows compilation preservation rates. Comment, preprocessor, and dead-branch carriers achieve perfect 100% compilation-preservation (1,434/1,434), confirming their non-interference at the compilation level; this is consistent with their construction (comments and preprocessor directives are stripped before compilation; dead-branch is never executed). Identifier Substitution achieves 90.9% (1,303/1,434). The 131 failures stem from collisions with injected header symbols (uint8_t, size_t, printf, malloc, etc.). These cases are counted as unsuccessful attacks in ASR computation (included in the $\mathrm{TP}_{\mathrm{clean}}$ denominator but not as flipped), providing a conservative ASR estimate.
@@ B.1 Dataset Construction Details
Data sources and target language. Unified-VUL-N contains vulnerable functions written in C/C++ and is constructed by pooling the vulnerable subsets from PrimeVul, BigVul, and DiverseVul. The goal is to evaluate robustness under syntax-preserving transformations on a single, fixed benchmark with controlled composition.
@@
Context budget and token-length filtering. To guarantee sufficient headroom for injected strings and to standardize the context budget across models, we filter functions by code-only token length using the Qwen2.5-Coder-14B tokenizer. Specifically, we require the code-only length to be at most 4096 tokens, where "code-only" refers to the function text used as input code (excluding any additional prompt or JSON schema wrappers). This constraint ensures that our injected variants remain within model context limits under the shared evaluation prompt.
@@ B.3 Benign-Class Invariance and False Positives
Setup. We evaluate benign-class invariance on a held-out set of $N_{\mathrm{benign}}{=}500$ ground-truth benign functions. Clean FPR is the fraction of benign functions predicted as VULNERABLE under clean inference (baseline model error). Sanitized FPR is computed after applying the same sanitization pipeline as in §5.3 (comment stripping and preprocessor removal); we report $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ as the change in false-positive rate (percentage points) from clean to sanitized inputs. Benign-class Invariance Rate is the fraction of benign functions that are predicted as BENIGN under clean inference and remain BENIGN after applying our syntax-preserving attacks (union over all attack methods). This metric directly measures whether our attacks preserve benign-class predictions, which is aligned with our attack objective (evading vulnerability detection without introducing false positives).
Results. Table 5 reports per-model clean FPR, sanitized FPR (and $\Delta\mathrm{FPR}_{\mathrm{sanit}}$), and benign-class invariance rate. All models achieve 100% benign-class invariance rate, demonstrating that our syntax-preserving attacks preserve benign-class predictions: all functions that are correctly classified as BENIGN under clean inference remain BENIGN after attack.
Table 5: Benign-side effects and invariance. Clean/Sanitized FPRs are measured on $N_{\mathrm{benign}}$ ground-truth benign functions. $\Delta\mathrm{FPR}_{\mathrm{sanit}}$ is the change (percentage points) after sanitization. Benign-class Invariance is computed on benign functions predicted as BENIGN under clean inference and remains BENIGN after applying the union of syntax-preserving attacks.
@@ B.4 Robustness of Reference Tokenizer for Length Filtering
To ensure fair cross-model comparisons, we standardized length filtering using the Qwen2.5-Coder-14B tokenizer to maintain an identical evaluation set across all targets. Per-model filtering would result in inconsistent subsets, making metrics like $\mathrm{TPR}_{\mathrm{clean}}$ and $\mathrm{ASR}_{\mathrm{cond}}$ non-comparable. Correlation analysis on 500 random samples confirms that the Qwen tokenizer is highly representative: it achieves a Pearson correlation of $r=0.9998$ with Llama-3.1-8B and $r=0.9941$ with Mistral-7B-Instruct-v0.2, which is the original model of CodeAstra. The mean length difference relative to Llama is only 5.57-5.57 tokens (max 16). Given our 4096-token limit, these negligible deviations ensure that the filtering process does not introduce systemic bias against specific architectures.
@@ B.5 Compute and Latency Statistics
Cost metrics. We track (i) forward passes for gradients ($m_{c}$ per step during curriculum, $m_{\mathrm{train}}$ per step after balance), (ii) forward passes for candidate evaluation ($m_{c}\cdot\lceil\texttt{search\_width}/\texttt{batch\_size}\rceil=m_{c}\cdot 32$ per step with our default settings), and (iii) wall-clock time (total optimization time including GPU memory transfers and batched operations). Optional per-step model.generate() calls are counted separately and can be disabled via enable_step_generate=False to reduce latency.
@@ C.1 Site Selection and Functional Role
Positional Sensitivity. Table 6 compares the $\mathrm{ASR}_{\mathrm{cond}}$ for different comment-carrier insertion sites. We observe that Trailing comments (placed after the function body) often exhibit high effectiveness, notably for Llama3.1-8B and CodeAstra, where success rates significantly exceed those of the original Leading header comments. However, Middle placements (interleaved with the logic) often show lower effectiveness across all models. This suggests that the model’s contextual focus on interleaved logic may partially mitigate the influence of syntax-preserving perturbations compared to injections at the function boundaries.
@@
Table 6: Ablation results for different injection sites and functional roles. $\mathrm{ASR}_{\mathrm{cond}}$ is reported as the fraction of originally correctly detected vulnerable functions ($\mathrm{TP}_{\mathrm{clean}}$) that were flipped to BENIGN; Qwen2.5-Coder-32B uses transfer strings; Llama3.1-8B and CodeAstra use on-target strings.
Model Family Variant Flipped / $\mathrm{TP}_{\mathrm{clean}}$ $\mathrm{ASR}_{\mathrm{cond}}$ (%)
@@
C.2 Transfer vs. On-Target Breakdown
Table 7 provides the full carrier-level comparison of transfer and on-target attack success across three white-box models. Table 8 shows surrogate-capacity ablation for Qwen2.5-Coder-32B identifier substitution.
Model Regime $\mathrm{ASR}_{\mathrm{cond}}$ (%)
Idsub Comment ifdef macro Dead-br
Llama3.1-8B Transfer 34.66 64.50 99.03 96.33 59.79
On-target 43.49 42.01 99.03 65.46 86.66
CodeAstra Transfer 4.72 78.60 42.59 75.48 72.12
On-target 85.41 63.06 16.62 98.87 97.51
StarCoder2-15B Transfer 1.31 4.00 0.18 86.83 1.58
On-target 16.27 69.59 5.45 85.25 32.86
On-target universal GCG (Qwen2.5-Coder-32B) 92.60 1038 / 1121
C.3 GCG Loss Trajectories
To support the hypothesis in §5.4 about carrier-level optimization difficulty, we ran GCG universal optimization on the Qwen2.5-Coder-14B surrogate for m=10m{=}10 vulnerable samples over 500 steps, recording per-step loss for both Identifier Substitution and Comment Insertion. After a curriculum phase completing at step 10, Identifier Substitution loss peaks at 2.89, then descends monotonically over ${\sim}300$ steps ($2.75\to 2.39\to 1.98\to 1.64$), stabilizing at 1.251.251.301.30. Comment Insertion, under the same protocol, spikes to 7.12 once all 10 samples are loaded, briefly drops to ${\sim}4.62$, then oscillates between $4.62$ and $4.75$ for the remaining 400 steps without further progress. These trajectories are consistent with Table 7: the constrained token space of valid C identifiers acts as a structural regularizer, aligning cross-sample gradients and enabling smooth convergence, whereas the unconstrained comment space allows gradients from diverse samples to interfere destructively, trapping the optimizer in a high-loss plateau.
Appendix D Extended Defense Analysis
[…]
@@ Appendix D Extended Defense Analysis
As shown in Table 10, the detection boundary is highly unstable. For Llama3.1-8B, simply removing developer comments (which have no effect on program execution) causes the model to "discover" 133 previously missed vulnerabilities—an 8.6% increase. This suggests that the model’s clean accuracy is partly contingent on specific lexical distributions rather than robust semantic understanding. Qwen2.5-Coder-32B shows similar behavior (+5.2%), while CodeAstra exhibits minimal drift (+0.09%), suggesting different sensitivities to non-executable context across architectures.